Find the angle between the vectors 𝑨⃗⃗ =(πŸ‘π’ŠΜ‚+πŸπ’‹Μ‚+π’ŒΜ‚) and 𝑨⃗⃗ =(πŸ“π’ŠΜ‚βˆ’πŸπ’‹Μ‚βˆ’πŸ‘π’ŒΜ‚)


1
Expert's answer
2021-03-08T08:32:17-0500

The angle between two vectors can be found as follows:


"A\cdot B=|A|\cdot|B|\cdot cos\theta,""cos\theta=\dfrac{A\cdot B}{|A|\cdot|B|},""cos\theta=\dfrac{A_x\cdot B_x+A_y\cdot B_y+A_z\cdot B_z}{\sqrt{A_x^2+A_y^2+A_z^2}\cdot\sqrt{B_x^2+B_y^2+B_z^2}},""\theta=cos^{-1}(\dfrac{A_x\cdot B_x+A_y\cdot B_y+A_z\cdot B_z}{\sqrt{A_x^2+A_y^2+A_z^2}\cdot\sqrt{B_x^2+B_y^2+B_z^2}}),""\theta=cos^{-1}(\dfrac{3\cdot5+2\cdot(-2)+1\cdot(-3)}{\sqrt{(3)^2+(2)^2+(1)^2}\cdot\sqrt{(5)^2+(-2)^2+(-3)^2}}),""\theta=69.7^{\circ}."